# Moduli stacks of Higgs bundles on stable curves

### v2

@@ 3.1.2 Space of expanded degenerations
$(\sigma_{1},\ldots,\sigma_{n})\cdot(t_{1},\ldots,t_{n+1}):=(\sigma_{1}\cdot t_{1},\ldots,\sigma_{i-1}^{-1}\cdot\sigma_{i}\cdot t_{i},\ldots,\sigma_{n}^{-1}\cdot t_{n+1}).$ (3.5)
@@
In [17, Section 1.1], Jun Li constructed a $G[n]$-equivariant family of expanded degenerations $W[n]$ over $S[n]$. It is constructed by several birational transformations of the family $\mathcal{X}\times_{S}S[n]$. The curves occurring in the family are all possible expanded degenerations of the family $\mathcal{X}/S$ for which the length of rational chains is bounded by $n$.
3.1.3 Stack of bounded Expanded degenerations
The significance of the family $W[n]\to S[n]$ is that it can be used as an atlas for the universal family of expanded degenerations with bubbling of length $\leq n$.
Indeed, recall from [17, Definition 1.9, Proposition 1.10] the following
@@ 3.1.3 Stack of bounded Expanded degenerations
Two such families $W_{T}$ and $W^{\prime}_{T}$ are isomorphic if there is a $T$-isomorphism $f:W_{T}\longrightarrow W^{\prime}_{S}$ compatible to the tautological projections $W_{T}\longrightarrow W$ and $W^{\prime}_{T}\longrightarrow W$.
@@
Here we use the following
@@
3.1.4 An alternative construction of the stack $\mathfrak{M}$
In [42, Definition 2.22], Zijun Zhou gave an alternative construction of the stack $\mathfrak{M}$ which will be useful for our purpose. Here we briefly recall his construction.
First let us introduce some useful notation and terminology. For any integer $n$, we set $[n]:=\{1,\dots,n+1\}$. It will be useful to have short names for the natural maps between the spaces $S[k]$ for various values of $k$.
@@ 3.1.4 An alternative construction of the stack $\mathfrak{M}$
Given any subset $I\subseteq[n]$ of cardinality $k+1$, we define
@@
    $\gamma_{I}:S[k]\hookrightarrow S[n]$ (3.7)
    given by $\gamma_{I}(t_{1},\dots,t_{k+1})=t_{i}$ if $i\in I$ and $\gamma_{I}(t_{1},\dots,t_{k+1})=1$ otherwise.
@@
    $\tau_{I}:S[k]\times\mathbb{G}_{m}^{n-k}\hookrightarrow S[n]$ (3.8)
@@
Given two subsets $I$ and $I^{\prime}$ of $[n]$ both of order $k+1$, define an equivalence relation $R_{I,I^{\prime}}$ on $S[n]$ by setting
$R_{I,I^{\prime}}=S[k]\times\mathbb{G}_{m}^{n-k}\rightrightarrows S[n]$ (3.9)
@@
$R_{k}:=\coprod_{|I|=|I^{\prime}|=k+1,I,I\subseteq[n]}R_{I,I^{\prime}}\ \rightrightarrows\ S[n].$ (3.10)
@@
$R_{dis}:=\coprod_{k\leq n}R_{k}\ \rightrightarrows\ S[n]$ (3.11)
Finally we define the total smooth groupoid
@@
From [42, Remark 2.23], it follows that the stack $\mathfrak{M}_{n}$ is the stack of expanded degenerations of the family $\mathcal{X}/S$ bounded by the integer $n$.
Notation:  From here onwards, we will only work with $\mathfrak{M}_{n}$ and we drop the subscript "$n"$ from the notation and denote it simply by $\mathfrak{M}$.
@@
$S[k]\times\mathbb{G}_{m}^{n-k}\rightrightarrows S[n]$ (3.13)
@@
For simplicity, let us assume that $S=\mathbb{A}^{1}$. Then $S[n]=\mathbb{A}^{n+1}$. Notice that $G[n]:=\mathbb{G}_{m}^{n}$. But in fact, there is an action of a bigger group $\mathbb{G}_{m}^{n+1}$ on $\mathbb{A}^{n+1}$ given by the coordinatewise multiplication. Moreover, the group $\mathbb{G}_{m}^{n}$ can be identified as a subgroup of $\mathbb{G}_{m}^{n+1}$ consisting of elements of the form $(t_{1},\dots,t_{n+1})$ such that $\prod^{n+1}_{i=1}t_{i}=1$. Let’s consider the action of the bigger group $G[n+1]$.
@@
Let us first define a natural action of $G[n+1]$ on $R_{I}:=\mathbb{A}^{k+1}\times\mathbb{G}_{m}^{n-k}$. Notice that the cardinality of the set $I$ equals $k+1$. It is an ordered subset of $[n+1]$. Therefore, the complement $J:=[n+1]\setminus I$ is also an ordered set of cardinality $n-k$, where the order is the induced order from $[n+1]$. Now we see that $\mathbb{G}_{m}^{n+1}\cong\mathbb{G}_{m}^{I}\times\mathbb{G}_{m}^{J}$. Finally, we define the action of $\mathbb{G}_{m}^{n+1}$ on $R_{I}:=\mathbb{A}^{I}\times\mathbb{G}_{m}^{J}$ by the obvious action of $\mathbb{G}_{m}^{I}\times\mathbb{G}_{m}^{J}$. It is easy to check that the map $R_{I}:\mathbb{A}^{k+1}\times(\mathbb{k}^{*})^{n-k}\longrightarrow\mathbb{A}^{n+1}$ is $\mathbb{G}_{m}^{n+1}$-equivariant. Hence, it is also equivariant under the action of the smaller subgroup $\mathbb{G}_{m}^{n}$. Therefore, the equivalence relations $R_{I,I^{\prime}}$ descends to the quotient stack $[S[n]/G[n]]$.
@@
Now we will show that the relation $\underline{R}_{dis}$ is an étale equivalence relation. First of all, notice that the maps $R_{I}:S[k]\times\mathbb{G}_{m}^{n-k}\longrightarrow S[n]$ are open immersions. Therefore, $R_{k}:=\coprod_{|I|=|I^{\prime}|=k+1,I,I\subseteq[n]}R_{I,I^{\prime}}\rightrightarrows\mathbb{A}^{k+1}$ is an étale equivalence relation because both the projections are Zariski open immersions. Therefore, $R_{dis}$ on $S[n]$ and $[R_{dis}/G[n]]$ on $[S[n]/G[n]]$ defines an étale equivalence relation. ∎
3.2 Construction of the family of curves
@@
In this subsection, we will briefly recall from [17, 42] the fact that there is a family of expanded degeneration (whose rational chain has length bounded by $n$). Let us choose an étale neighborhood $U_{p}$ of the node $p\in\mathcal{X}$ such that the nodal curve $U_{p}\times_{\mathcal{X}}X_{0}$ is a reducible nodal curve with two smooth connected components intersecting transversally at the node $p$. Let $V=\mathcal{X}-\{p\}$. Then $\{U_{p},V\}$ is an étale covering of $\mathcal{X}$. We have the following diagram (which is both cartesian and co-cartesian/push-out):
${U_{p}\times_{\mathcal{X}}V}$${V}$${U_{p}}$${\mathcal{X}}$ (3.14)
@@
Notice that $U_{p}\longrightarrow S$ is a simple normal crossing degeneration which we can feed into the construction of [17, 42] to produce a family of expanded degenerations. We denote this family by $U_{p}[n]\longrightarrow S[n]$ and also write $V[n]:=V\times_{S}S[n]$. Now we construct the family of expanded degenerations, denoted by $\mathcal{X}[n]$ over $S[n]$ by the following push-out
${U_{p}[n]\times_{(\mathcal{X}\times_{S}S[n])}V[n]}$${V[n]}$${U_{p}[n]}$${\mathcal{X}[n]}$ (3.15)
@@
The push-out exists as an algebraic space by Rydh’s pushout theorem [34, Theorem C]. The family of curves $\mathcal{X}[n]\longrightarrow S[n]$ is the desired family of expanded degenerations of the original family $\mathcal{X}\longrightarrow S$. We refer to [42, Proposition 2.13] for the key properties of the family of expanded degenerations. In particular, by properties (3) and (4) of [42, Proposition 2.13] the family $U_{p}[n]\to S[n]$ is equivariant under the action of the smooth groupoid $R_{tot}\rightrightarrows S[n]$. By the universality of pushouts this implies that the family of expanded degenerations $\mathcal{X}[n]\to S[n]$ is also equivariant and so $\mathcal{X}[n]$ descends to the stack of expanded degenerations $\mathfrak{M}=\mathfrak{M}_{n}=\left[S[n]/R_{tot}\right]$. We denote this family of expanded degenerations by $\mathfrak{X}_{\mathfrak{M}}\to\mathfrak{M}$.
3.3 Log structures on derived Artin stacks
@@
In this subsection, we discuss shifted log-symplectic structure on a log quasi-smooth derived Artin stack (Definition 3.4.0(b)) equipped with a locally free log structure. We refer to [26, 27] and [29] for the prerequisite material on log-structure/locally free log-structure on Artin/derived Artin stacks. We now recall a few necessary definitions.
@@
Let $\mathfrak{L}^{0}$ denote the algebraic stack which classifies fine log structures, and let $\mathfrak{L}^{1}$ denote the algebraic stack which classifies morphisms of fine log schemes [27]. We denote by $\mathfrak{L}^{0}_{\mathsf{lf}}$ and $\mathfrak{L}^{1}_{\mathsf{lf}}$ the classifying substacks (known to be smooth by [29, Proposition 1.7]) of locally free log structures and morphisms between locally free log structures, respectively. Given a logarithmic scheme $(S,P,\alpha)$, the stack $\mathfrak{L}^{0}_{S,\mathsf{lf}}:=\mathfrak{L}^{1}_{\mathsf{lf}}\times_{\mathfrak{L}^{0}_{\mathsf{lf}}}S$ is the classifying stack of logarithmic morphisms from a scheme with locally free log structure to $(S,P,\alpha)$. With this notation, we now have the following natural definition [29]:
@@
${\mathfrak{X}}$${\mathfrak{L}^{1}_{\mathsf{lf}}}$${\mathfrak{Y}}$${\mathfrak{L}^{0}_{\mathsf{lf}}}$ (3.16)
3.3.1 Relative logarithmic cotangent complex
@@
${\mathbb{L}}^{log}_{\mathfrak{X}/\mathfrak{Y}}:=\mathbb{L}_{\mathfrak{X}/\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}}$ (3.17)
Here, as before, $\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}:=\mathfrak{L}^{1}_{\mathsf{lf}}\times_{\mathfrak{Y}}\mathfrak{L}^{0}_{\mathsf{lf}}$ and $\mathbb{L}_{\mathfrak{X}/\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}}$ denotes the relative cotangent complex for the morphism of stacks $\mathfrak{X}\longrightarrow\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}$.
With this definition in mind, we return to the discussion on the stack of bounded expanded degenerations $\mathfrak{M}\rightarrow S$ (3.4, 3.8) for a given one parameter degeneration $\mathcal{X}\rightarrow S$ of a smooth projective curve to a nodal curve.
@@
The morphism $\mathfrak{M}\longrightarrow S$ of ordinary Artin stacks is a log-smooth map. Moreover, the relative log-cotangent complex $\mathbb{L}^{log}_{\mathfrak{M}/S}=0$.
@@
From Lemma 3.9, it follows that the map $[S[n]/G[n]]\longrightarrow\mathfrak{M}$ is étale. Notice that the varieties $S[n]$ and $S$ are log smooth varieties with the log structure coming from the divisor given by the pre-image of the closed point of $S$ under the map $S[n]\longrightarrow S$. Since the map $S[n]\longrightarrow S$ is an étale base-change of the map $\mathbb{A}^{n+1}\longrightarrow\mathbb{A}^{1}$ given by $(t_{1},\dots,t_{n+1})\mapsto t_{1}\cdots t_{n+1}$, the map is a log-smooth morphism. Therefore, it follows that the map $[S[n]/G[n]]\longrightarrow S$ is also a log-smooth morphism of log-smooth Artin stacks.
The explicit description of the relative log-cotangent complex is the following. The pull-back (to $S[n]$) of the log-cotangent complex of the stack $[S[n]/G[n]]$ is the perfect complex on $S[n]$ concentrated in degrees 00 and $1$ which is given by
$\left[\Omega_{S[n]}^{1}(\log\leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}\right]$ (3.18)
where $\Omega_{S[n]}^{1}(\log\leavevmode\nobreak\ \partial S[n])$ is the rank $(n+1)$ locally free sheaf of logarithmic one forms on $S[n]$ with poles along the strict normal crossings divisor $\partial S[n]=S[n]\times_{S}\{o\}$. Since the action of $G[n]$ on $S[n]$ respects $\partial S[n]$, we get a natural $G[n]$-equivariant structure $\Omega_{S[n]}^{1}(\log\leavevmode\nobreak\ \partial S[n])$ such that the natural log-symplectic structure is $G[n]$-equivariant. The morphism (3.18) is given by the (equivariant) moment map for the action of $G[n]$ on the log-cotangent bundle of $S[n]$. In particular (3.18) is a $G[n]$-perfect complex. We recall that the action of $G[n]$ on $S[n]$ is given by
$(t_{1},\dots,t_{n+1})\cdot(x_{1},\dots,x_{n+1})=(t_{1}x_{1},t_{2}x_{2},\dots,t_{n+1}x_{n+1})$ (3.19)
where $t_{1}\cdots t_{n+1}=1$, and that the map $S[n]\longrightarrow S$ is given by
$(t_{1},\dots,t_{n+1})\mapsto\prod^{n+1}_{i=1}t_{i}$ (3.20)
It is easy to see that the map 3.18
$\left[\Omega_{S[n]}^{1}(\log\leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}\right]$ (3.21)
of vector bundles is surjective: consider the action of $(\mathbb{G}_{m})^{n}$ on $\mathbb{A}^{n+1}$ given by 3.19. Let $p:=(x_{1},\cdots,x_{n+1})\in\mathbb{A}^{n+1}$ be any point. Then consider the orbit map at the point $p$.
$Or_{p}:(\mathbb{G}_{m})^{n}\longrightarrow\mathbb{A}^{n+1}$ (3.22)
$(t_{1},\cdots,t_{n+1})\mapsto(t_{1}\cdot x_{1},\cdots,t_{n+1}\cdot x_{n+1})$ (3.23)
The induced map on the tangent spaces is $d(Or)_{p}:T_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}=\mathbb{C}^{n}\longrightarrow T_{p}\mathbb{A}^{n+1}=\mathbb{C}^{n+1}$ is given by the multiplication by the following diagonal matrix

### v3

@@ 3.1.2. Space of expanded degenerations
(3.5) $(\sigma_{1},\ldots,\sigma_{n})\cdot(t_{1},\ldots,t_{n+1}):=(\sigma_{1}\cdot t_{1},\ldots,\sigma_{i-1}^{-1}\cdot\sigma_{i}\cdot t_{i},\ldots,\sigma_{n}^{-1}\cdot t_{n+1}).$
@@
In [21, Section 1.1], Jun Li constructed a $G[n]$-equivariant family of expanded degenerations $W[n]$ over $S[n]$. It is constructed by several birational transformations of the family $\mathcal{X}\times_{S}S[n]$. The curves that occur in the family are all possible expanded degenerations of the family $\mathcal{X}/S$ for which the length of rational chains is bounded by $n$.
Notation: To be consistent with our notations, from now on, we will denote Jun Li’s family of expanded degenerations by $\mathfrak{X}[n]$ over $S[n]$.
3.1.3. Stack of bounded expanded degenerations
The significance of the family $\mathfrak{X}[n]\to S[n]$ is that it can be used as an atlas for the universal family of expanded degenerations with bubbling of length $\leq n$.
Indeed, recall from [21, Definition 1.9, Proposition 1.10] the following
@@ 3.1.3. Stack of bounded expanded degenerations
Two such families $\mathfrak{X}_{T}$ and $\mathfrak{X}^{\prime}_{T}$ are isomorphic if there is a $T$-isomorphism $f:\mathfrak{X}_{T}\longrightarrow\mathfrak{X}^{\prime}_{T}$ compatible with tautological projections $\mathfrak{X}_{T}\longrightarrow\mathcal{X}$ and $\mathfrak{X}^{\prime}_{T}\longrightarrow\mathcal{X}$.
Remark 3.7.
@@
The following properties are obtained from the construction / Definition of the expanded degeneration stack as the quotient of $\mathbb{A}^{n+1}$ by a smooth equivalence relation (see [44, page 16, Def. 2.22.]).
Here we use the following.
@@
3.1.4. An alternative construction of the stack $\mathfrak{M}$
In [44, Definition 2.22], Zijun Zhou gave an alternative construction of the stack $\mathfrak{M}$ which will be useful for our purposes. Here we briefly recall his construction.
First, let us introduce some useful terminology and notation. For any integer $n$, we set $n]:=\{1,\dots,n+1\}$. It will be useful to have short names for the natural maps between the spaces $S[k]$ for various values of $k$.
@@ 3.1.4. An alternative construction of the stack $\mathfrak{M}$
Given any subset $I\subseteq[n]$ of cardinality, say $k+1$, we define
@@
    (3.7) $\gamma_{I}:S[k]\hookrightarrow S[n]$
    given by $\gamma_{I}(t_{1},\dots,t_{k+1})=(z_{1},\dots,z_{n+1})$, where
    $(z_{1},\dots,z_{n+1})=\begin{cases}t_{i},&\text{when}\penalty 10000\ \penalty 10000\ i\in I\\ 1,&\text{when}\penalty 10000\ \penalty 10000\ i\in[n]\setminus I\end{cases}$
    Notice that $S[k]:=S\times_{\mathbb{A}^{1}}\mathbb{A}^{k+1}$ and $(t_{1},\dots,t_{k+1})$ denote a point of $S[k]$.
@@
    (3.8) $\tau_{I}:S[k]\times\mathbb{G}_{m}^{n-k}\hookrightarrow S[n]$
@@
(i):
Given two subsets $I$ and $I^{\prime}$ of $[n]$ both of order $k+1$, define the equivalence relation $R_{I,I^{\prime}}$ on $S[n]$ by setting
(3.9) $R_{I,I^{\prime}}=S[k]\times\mathbb{G}_{m}^{n-k}\rightrightarrows S[n]$
@@
(3.10) $R_{k}:=\coprod_{|I|=|I^{\prime}|=k+1,I,I^{\prime}\subseteq[n]}R_{I,I^{\prime}}\ \rightrightarrows\ S[n].$
(iii):
@@
(3.11) $R_{dis}:=\coprod_{k\leq n}R_{k}\ \rightrightarrows\ S[n]$
(iv):
Finally, we define the total smooth groupoid
@@
From [44, Remark 2.23], it follows that the stack $\mathfrak{M}_{n}$ is the stack of expanded degenerations of the family $\mathcal{X}/S$ bounded by the integer $n$.
Notation: From here onwards, we will only work with $\mathfrak{M}_{n}$ and we drop the subscript $n"$ from the notation and denote it simply by $\mathfrak{M}$.
@@
(3.13) $S[k]\times\mathbb{G}_{m}^{n-k}\rightrightarrows S[n]$
@@
For simplicity, assume that $S=\mathbb{A}^{1}$. Then $S[n]=\mathbb{A}^{n+1}$. Notice that $G[n]:=\mathbb{G}_{m}^{n}$. But in fact, there is an action of a larger group $\mathbb{G}_{m}^{n+1}$ on $\mathbb{A}^{n+1}$ given by coordinate multiplication. Moreover, the group $\mathbb{G}_{m}^{n}$ can be identified as a subgroup of $\mathbb{G}_{m}^{n+1}$ consisting of elements of the form $(t_{1},\dots,t_{n+1})$ such that $\prod^{n+1}_{i=1}t_{i}=1$. Let us consider the action of the larger group $G[n+1]$.
@@
Let us first define the natural action of $G[n+1]$ on $R_{I}:=\mathbb{A}^{k+1}\times\mathbb{G}_{m}^{n-k}$. Notice that the cardinality of the set $I$ is equal to $k+1$. It is an ordered subset of $[n+1]$. Therefore, the complement $J:=[n+1]\setminus I$ is also an ordered set of cardinality nkn-k, where the order is the induced order of $[n+1]$. Now we see that $\mathbb{G}_{m}^{n+1}\cong\mathbb{G}_{m}^{I}\times\mathbb{G}_{m}^{J}$. Finally, we define the action of $\mathbb{G}_{m}^{n+1}$ on $R_{I}:=\mathbb{A}^{I}\times\mathbb{G}_{m}^{J}$ by the obvious action of $\mathbb{G}_{m}^{I}\times\mathbb{G}_{m}^{J}$. It is easy to check that the map $R_{I}:\mathbb{A}^{k+1}\times(\mathbb{k}^{*})^{n-k}\longrightarrow\mathbb{A}^{n+1}$ is $\mathbb{G}_{m}^{n+1}$-equivariant. Hence, it is also equivariant under the action of the smaller subgroup $\mathbb{G}_{m}^{n}$. Therefore, the equivalence relations $R_{I,I^{\prime}}$ descend to the quotient stack $[S[n]/G[n]]$.
@@
Now we will show that the relation $\underline{R}_{dis}$ is an étale equivalence relation. First of all, note that the maps $R_{I}:S[k]\times\mathbb{G}_{m}^{n-k}\longrightarrow S[n]$ are open immersions. Therefore, $R_{k}:=\coprod_{|I|=|I^{\prime}|=k+1,I,I\subseteq[n]}R_{I,I^{\prime}}\rightrightarrows\mathbb{A}^{k+1}$ is an étale equivalence relation because both the projections are Zariski open immersions. Therefore, $R_{dis}$ on $S[n]$ and $R_{dis}/G[n]]$ on $[S[n]/G[n]]$ define an equivalence relation étale equivalence relation. ∎
3.2. Construction of the family of curves
@@
In this subsection, we will briefly recall from [21, 44] the fact that there is a family of expanded degeneration (whose rational chain has a length bounded by $n$). Let us choose a neighborhood étale $U_{p}$ of node $p\in\mathcal{X}$ such that the nodal curve $U_{p}\times_{\mathcal{X}}X_{0}$ is a reducible nodal curve with two smooth connected components intersecting transversely at node $p$. Let $V=\mathcal{X}-\{p\}$. Then $\{U_{p},V\}$ is an étale covering of $\mathcal{X}$. We have the following diagram (which is both cartesian and co-cartesian/push-out):
[…]
@@
Notice that $U_{p}\longrightarrow S$ is a simple normal crossing degeneration that we can feed into the construction of [21, 44] to produce a family of expanded degenerations. We denote this family by $U_{p}[n]\longrightarrow S[n]$ and also write $V[n]:=V\times_{S}S[n]$. Now we construct the family of expanded degenerations, denoted by $\mathcal{X}[n]$ over $S[n]$, by the following push-out
[…]
The pushout exists as an algebraic space according to Rydh’s pushout theorem [34, Theorem C]. The family of curves $\mathcal{X}[n]\longrightarrow S[n]$ is the desired family of expanded degenerations of the original family $\mathcal{X}\longrightarrow S$.
@@
We refer to [44, Proposition 2.13] for the key properties of the expanded degeneration family. In particular, by properties (3) and (4) of [44, Proposition 2.13] the family $U_{p}[n]\to S[n]$ is equivariant under the action of the smooth groupoid $R_{tot}\rightrightarrows S[n]$. By the universality of pushouts, this implies that the family of expanded degenerations $\mathcal{X}[n]\to S[n]$ is also equivariant and so $\mathcal{X}[n]$ descends to the stack of expanded degenerations $\mathfrak{M}=\mathfrak{M}_{n}=\left[S[n]/R_{tot}\right]$. We denote this family of expanded degenerations by $\mathfrak{X}_{\mathfrak{M}}\to\mathfrak{M}$.
Proposition 3.13.
The family $\mathfrak{X}_{\mathfrak{M}}\to\mathfrak{M}$ has the following universal property: Given any family of expanded degenerations (bounded by the integer $n$) $\mathfrak{X}_{T}\to T$ (over a scheme $T$) of the family of curves $\mathcal{X}/S$, there exists a unique morphism $\rho:T\to\mathfrak{M}_{n}$ such that $\rho^{*}\mathcal{X}\cong\mathfrak{X}_{T}$.
Proof.
The proof follows from [21, Lemma 1.8. , Definition 1.9., Proposition 1.10.] and [44, 2.3 Standard families of expanded degenerations and pairs – gluing, and 2.5 Stacks of expanded degenerations and pairs]. The Lemma [21, Lemma 1.8.] proves the universal property of the stack and the expanded family for the case when $\mathcal{X}/S$ is a ”simple-degeneration”, which means that the closed fiber of the family $\mathcal{X}/S$ is a union of two smooth curves that intersect transversely at a point. Then following the arguments of [44, 2.3 Standard families of expanded degenerations and pairs – gluing, and 2.5 Stacks of expanded degenerations and pairs] the universal property follows in our case, i.e., when the closed fibre is an irreducible nodal curve. Notice that even though Zijun explained the construction for the simple-degeneration case (for higher dimensional DM-stacks), his arguments work for our case as well, which is the case when the closed fiber of $\mathcal{X}/S$ is an irreducible nodal curve. We merely explained the construction here because for an irreducible nodal curve the construction involves a much simpler choice of etale covers. ∎
3.3. Log structures on derived Artin stacks
@@
In this subsection, we discuss locally-free log structures on a quasi-smooth derived Artin stack. We refer to [27, 28] and [33] for the prerequisite material on log-structure/ locally free log-structure on Artin/ derived Artin stacks. We now recall a few necessary definitions.
@@
Let $\mathfrak{L}^{0}$ denote the algebraic stack that classifies fine log structures, and let $\mathfrak{L}^{1}$ denote the algebraic stack that classifies morphisms of fine log schemes [28]. We denote by $\mathfrak{L}^{0}_{\mathsf{lf}}$ and $\mathfrak{L}^{1}_{\mathsf{lf}}$ the classifying substacks (known to be smooth by [33, Proposition 1.7]) of locally free log structures and morphisms between locally free log structures, respectively. Given a logarithmic scheme $(S,P,\alpha)$, the stack $\mathfrak{L}^{0}_{S,\mathsf{lf}}:=\mathfrak{L}^{1}_{\mathsf{lf}}\times_{\mathfrak{L}^{0}_{\mathsf{lf}}}S$ is the classifying stack of logarithmic morphisms from a scheme with locally free log structure to $(S,P,\alpha)$. With this notation, we now have the following natural definition [33]:
@@
(3.16) $\mathfrak{X}$$\mathfrak{L}^{1}_{\mathsf{lf}}$$\mathfrak{Y}$$\mathfrak{L}^{0}_{\mathsf{lf}}$}
3.3.1. Relative logarithmic cotangent complex
@@
(3.17) ${\mathbb{L}}^{\log}_{\mathfrak{X}/\mathfrak{Y}}:=\mathbb{L}_{\mathfrak{X}/\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}}$
Here, as before, $\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}:=\mathfrak{L}^{1}_{\mathsf{lf}}\times_{\mathfrak{L}^{0}_{\mathsf{lf}}}\mathfrak{Y}$ and $\mathbb{L}_{\mathfrak{X}/\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}}$ denote the relative cotangent complex for the morphism of stacks $\mathfrak{X}\longrightarrow\mathfrak{L}^{0}_{\mathfrak{Y},\mathsf{lf}}$.
With this definition in mind, we return to the discussion on the stack of bounded expanded degenerations $\mathfrak{M}\rightarrow S$ (3.6, 3.11) for a given one parameter degeneration $\mathcal{X}\rightarrow S$ of curves.
@@
The morphism $\mathfrak{M}\longrightarrow S$ is a log-smooth map. Moreover, the relative log-cotangent complex $\mathbb{L}^{\log}_{\mathfrak{M}/S}=0$.
@@
From Lemma 3.12, it follows that the map $S[n]/G[n]]\longrightarrow\mathfrak{M}$ is étale. Notice that the varieties $S[n]$ and $S$ are log smooth varieties with the log structure coming from the divisor given by the pre-image of the closed point of $S$ under the map $S[n]\longrightarrow S$. Since the map $S[n]\longrightarrow S$ is an étale base-change of the map $\mathbb{A}^{n+1}\longrightarrow\mathbb{A}^{1}$ given by $(t_{1},\dots,t_{n+1})\mapsto t_{1}\cdots t_{n+1}$, the map is a log-smooth morphism. Therefore, it follows that the map $S[n]/G[n]]\longrightarrow S$ is also a log-smooth morphism of log-smooth Artin stacks.
The explicit description of the relative log-cotangent complex is the following. The pull-back (to $S[n]$) of the log-cotangent complex of the stack $[S[n]/G[n]]$ is the perfect complex on $S[n]$ concentrated in degrees 0 and $1$ which is given by
(3.18) $\left[\Omega_{S[n]}^{1}(\log\penalty 10000\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}\right]$
where $\Omega_{S[n]}^{1}(\log\penalty 10000\ \partial S[n])$ is the rank $(n+1)$ locally free sheaf of logarithmic one forms on $S[n]$ with poles along the strict normal crossings divisor $\partial S[n]=S[n]\times_{S}\{o\}$. Since the action of $G[n]$ on $S[n]$ respects $\partial S[n]$, we obtain a natural $G[n]$-equivariant structure $\Omega_{S[n]}^{1}(\log\penalty 10000\ \partial S[n])$ such that the natural log-symplectic structure is $G[n]$-equivariant. The morphism (3.18) is given by the (equivariant) moment map for the action of $G[n]$ on the log-cotangent bundle of $S[n]$. In particular (3.18) is a $G[n]$-perfect complex. We recall that the action of $G[n]$ on $S[n]$ is given by
(3.19) $(t_{1},\dots,t_{n+1})\cdot(x_{1},\dots,x_{n+1})=(t_{1}x_{1},t_{2}x_{2},\dots,t_{n+1}x_{n+1})$
where $t_{1}\cdots t_{n+1}=1$, and the map $S[n]\longrightarrow S$ is given by
(3.20) $(t_{1},\dots,t_{n+1})\mapsto\prod^{n+1}_{i=1}t_{i}$
It is easy to see that the map (3.18)
(3.21) $\left[\Omega_{S[n]}^{1}(\log\penalty 10000\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}\right]$
of vector bundles is surjective: consider the action of $(\mathbb{G}_{m})^{n}$ on $\mathbb{A}^{n+1}$ given by (3.19). Let $p:=(x_{1},\cdots,x_{n+1})\in\mathbb{A}^{n+1}$ be any point. Then consider the orbit map at the point $p$.
(3.22) $Or_{p}:(\mathbb{G}_{m})^{n}\longrightarrow\mathbb{A}^{n+1}$
(3.23) $(t_{1},\cdots,t_{n+1})\mapsto(t_{1}\cdot x_{1},\cdots,t_{n+1}\cdot x_{n+1})$
The induced map on the tangent spaces is $d(Or)_{p}:T_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}=\mathbb{k}^{n}\longrightarrow T_{p}\mathbb{A}^{n+1}=\mathbb{k}^{n+1}$ is given by the multiplication by the following diagonal matrix